Find dy/dx for (1) y = t^2, x = 2t^3 (2) y = log(t^2), t^2 = x^3 (3) y = sin x , x = cos x

Find dy/dx for
(1) y = t^2, x = 2t^3
(2) y = log(t^2), t^2 = x^3
(3) y = sin x , x = cos x

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
I am solving 1st part-
dy/dx = dy/dt / dx/dt
= 2t/6t²
= 1/3t
Now
x = 2 t³
x = 2 t *y
t = x/2y
Hence
dy/dx = 2y/3x

Regards
Arun (askIITians forum expert)